Solve the problem.

Annual precipitation in a certain city is normally distributed with a mean of 99 inches, and a standard deviation of 18 in. Find the probability that the mean annual precipitation during 35 randomly picked years will be less than 101.8 in.?

Group of answer choices

0.8212

0.3212

0.9203

0.6788

0.1788
Answer #1

µ =    99      
σ =    18      
n=   35      
          
X =   101.8      
Z =(X - µ )/(σ/√n) =    (101.8-99)/(18/√35)=   0.92  
P(X<101.8) = P(Z<0.92) =        0.8212 (answer)